# Learning Mean-Field Games through Mean-Field Actor-Critic Flow

### v1

@@ B.3 Superlinear growth lemma
We prove (B.10) in this section. The motivation comes from Lemma A.9, which shows that the difference between the value function and the optimal value function has a superlinear (quadratic in (A.18)) growth with respect to the difference in controls.
@@
Assume $\mu\in\mathcal{M}$, $\alpha,\alpha^{\mu,*}\in\mathcal{A}$, $\chi=\frac{2}{d+5}$. Then,
$\left\|\nabla_{x}V^{\mu,\alpha}-\nabla_{x}V^{\mu,*}\right\|_{\mu,\alpha}\leq C\left\|\alpha-\alpha^{\mu,*}\right\|_{\mu,\alpha}^{1+\chi}.$ (B.95)
@@
Let $(t,x)\in[0,T]\times\mathbb{R}^{d}$. Let $x_{s}=X^{\mu,\alpha}_{s}$ be the state dynamic under $\mu$ and $\alpha$, with initialization $x_{t}=x$. Let $\alpha_{s}=\alpha(x,x_{s})$, $\alpha^{*}_{s}=\alpha^{\mu,*}(s,x_{s})$. We denote $\phi:=\alpha-\alpha^{\mu,*}$ and $\phi_{s}:=\phi(s,x_{s})$. By (A.29) in Lemma A.13, we only need to show (B.2) when $\left\|\phi\right\|_{\mu,\alpha}\leq 1$. By Lemma A.9,
@@
Step 1. We want to bound $\left\|\partial_{x_{1}}V^{\mu,\alpha}-\partial_{x_{1}}V^{\mu,*}\right\|_{\mu,\alpha}$. Then we can apply the same analysis in each dimension and get a bound for $\left\|\nabla_{x}V^{\mu,\alpha}-\nabla_{x}V^{\mu,*}\right\|_{\mu,\alpha}$. Given $(t,x)$, we define the tangent SDE [37] as $y_{s}=\partial_{x_{1}}X_{s}^{\mu,\alpha}$, i.e., the partial derivative of $x_{s}=X^{\mu,\alpha}_{s}$ with respect to the first argument of its initial condition. We denote $b^{\mu,\alpha}(t,x):=b(t,x,\mu_{t},\alpha(t,x))$ and $\sigma^{\mu}(t,x):=\sigma(t,x,\mu_{t})$. Then $y_{s}$ satisfies $y_{t}=e_{1}$ and
$\mathrm{d}y_{s}=\nabla_{x}b^{\mu,\alpha}(t,x_{s})y_{s}\,\mathrm{d}s+\left(\nabla_{x}\sigma^{\mu}(s,x_{s})\cdot y_{s}\right)\mathrm{d}W_{s}$
where $\nabla_{x}\sigma^{\mu}(s,x_{s})\cdot y_{s}:=\sum_{i=1}^{d}\partial_{x_{i}}\sigma^{\mu}(s,x)(y_{s})_{i}$. We denote $\sigma^{y}_{s}:=\nabla_{x}\sigma^{\mu}(s,x_{s})\cdot y_{s}$ for simplicity. By Ito’s formula,
@@
Since $\left|\nabla_{x}b^{\mu,\alpha}\right|=\left|\nabla_{x}b+\nabla_{\alpha}b\,\nabla_{x}\alpha\right|$ and $\left|\sigma^{\mu}\right|$ are bounded, we have
@@
Together with $|y_{t}|=1$, we know from Grönwall’s inequality that
$\mathbb{E}\left[\left|y_{s}\right|^{2}~\big|~x_{t}=x\right]\leq C$ (B.97)
where this constant $C$ is independent of $t$ and $x$.
Next, we want to estimate $\partial_{x_{1}}V^{\mu,\alpha}-\partial_{x_{1}}V^{\mu,*}$ using (B.96) and $y_{s}$. We recall that
$\nabla_{\alpha}^{2}H\left(s,x_{s},\mu_{s},(\alpha^{*}_{s}+v\phi_{s})(s,x_{s}),-\nabla_{x}V^{\mu,*}(s,x_{s})\right)=-\nabla_{\alpha}^{2}f-\sum_{i=1}^{d}\nabla_{\alpha}^{2}b_{i}\,\partial_{x_{i}}V^{\mu,*}(s,x_{s})$
Taking derivative with respect to $(x_{t})_{1}$ (first with respect to $x_{s}$, then multiply $\partial_{(x_{t})_{1}}x_{s}=y_{s}$ from chain rule), we obtain
@@
where we omit the input $(s,x_{s},\mu_{s},(\alpha^{*}+v\phi)(s,x_{s}))$ for $f$ and $b_{i}$ (and their derivatives); and the input $(s,x_{s})$ for $\alpha^{*}+v\phi$ and $V^{\mu,*}$ (and their derivatives). Using the bounds in Assumption 4.1, the bounds for $\alpha$, $\alpha^{\mu,*}$ and their gradients, and the estimation for the value function in Lemma A.12, we obtain
$\left|\partial_{(x_{t})_{1}}\nabla_{\alpha}^{2}H\left(s,x_{s},\mu_{s},(\alpha^{*}_{s}+v\phi_{s})(s,x_{s}),-\nabla_{x}V^{\mu,*}(s,x_{s})\right)\right|\leq C(1+|x_{s}|)|y_{s}|$ (B.98)
and
@@
For the term $\phi(s,x_{s})$, we have $\partial_{(x_{t})_{1}}\phi(s,x_{s})=\nabla_{x}\phi(s,x_{s})y_{s}$. Therefore, taking derivative of (B.96) with respect to $x_{1}$, we get
$\partial_{x_{1}}V^{\mu,\alpha}(t,x)-\partial_{x_{1}}V^{\mu,\alpha}(t,x)=-\mathbb{E}\Big[\int_{0}^{T}\int_{0}^{1}\int_{0}^{u}\Big(2\phi_{s}^{\top}\nabla_{\alpha}^{2}H\,\nabla_{x}\phi_{s}\,y_{s}+\phi_{s}^{\top}(\partial_{(x_{t})_{1}}\nabla_{\alpha}^{2}H)\phi_{s}\Big)\mathrm{d}v\,\mathrm{d}u\,\mathrm{d}s\Big]$ (B.100)
where $\nabla_{x}\phi_{s}:=\nabla_{x}\phi(s,x_{s})$ and we omit the input $\left(s,x_{s},\mu_{s},(\alpha^{*}_{s}+v\phi_{s})(s,x_{s}),-\nabla_{x}V^{\mu,*}(s,x_{s})\right)$ for $\nabla_{\alpha}^{2}H$.
Substitute the estimations (B.98), (B.99) into (B.100), we get
$\quad\left|\partial_{x_{1}}V^{\mu,\alpha}(t,x)-\partial_{x_{1}}V^{\mu,\alpha}(t,x)\right|$ (B.101)
$\leq C\,\mathbb{E}\Big[\int_{t}^{T}\int_{0}^{1}\int_{0}^{u}(1+|x_{s}|)|\phi_{s}|\left(|\phi_{s}|+|\nabla_{x}\phi_{s}|\right)|y_{s}|\,\mathrm{d}v\,\mathrm{d}u\,\mathrm{d}s~\Big|~x_{t}=x\Big]$
$\leq C\,\mathbb{E}\Big[\int_{t}^{T}(1+|x_{s}|)|\phi_{s}|\left(|\phi_{s}|+|\nabla_{x}\phi_{s}|\right)|y_{s}|\,\mathrm{d}s~\Big|~x_{t}=x\Big]$
Therefore
$\quad\left\|\partial_{x_{1}}V^{\mu,\alpha}-\partial_{x_{1}}V^{\mu,\alpha}\right\|^{2}_{\mu,\alpha}$
$\leq C\,\mathbb{E}\Big\{\int_{0}^{T}\mathbb{E}\Big[\int_{t}^{T}(1+|x_{s}|)|\phi_{s}|\left(|\phi_{s}|+|\nabla_{x}\phi_{s}|\right)|y_{s}|\,\mathrm{d}s~\Big|~x_{t}=x\Big]^{2}\mathrm{d}t\Big\}$
$\leq C\,\mathbb{E}\Big\{\int_{0}^{T}\mathbb{E}\Big[\int_{t}^{T}(1+|x_{s}|)^{2}|\phi_{s}|^{2}\left(|\phi_{s}|+|\nabla_{x}\phi_{s}|\right)^{2}\,\mathrm{d}s~\Big|~x_{t}=x\Big]\cdot\Big[\int_{t}^{T}|y_{s}|^{2}\,\mathrm{d}s~\Big|~x_{t}=x\Big]\,\mathrm{d}t\Big\}$
$\leq C\,\mathbb{E}\Big\{\int_{0}^{T}\mathbb{E}\Big[\int_{t}^{T}(1+|x_{s}|^{2})|\phi_{s}|^{2}\left(|\phi_{s}|+|\nabla_{x}\phi_{s}|\right)^{2}\,\mathrm{d}s~\Big|~x_{t}=x\Big]\,\mathrm{d}t\Big\}$
@@
We consecutively used: estimation (B.101), Hölder’s inequality, estimation (B.97), (exchanging the integration in $s$ and $t$ and) tower property in the four inequalities above. Applying the same analysis in each dimension, we obtain
@@
Step 2. We estimate the right-hand side of (B.102) in this step. Recall the density function $\rho$ satisfies the Aronson-type bound (4.1). We fix $R>0$ s.t. $1+R^{2}\geq 2K$. We denote $B_{R}:=\{x\in\mathbb{R}^{d}:|x|\leq R\}$ and $B_{R}^{c}:=\{x\in\mathbb{R}^{d}:|x|>R\}$. Since (omit the input $(t,x)$ for $\phi$, $\Phi$, and $\rho$ temporally)
$(1+R^{2})\int_{0}^{T}\int_{B_{R}^{c}}\left|\Phi\right|^{2}\rho\,\mathrm{d}x\,\mathrm{d}t\leq\int_{0}^{T}\int_{\mathbb{R}^{d}}(1+|x|^{2})\left|\Phi\right|^{2}\rho\,\mathrm{d}x\,\mathrm{d}t\leq K\int_{0}^{T}\int_{\mathbb{R}^{d}}\left|\Phi\right|^{2}\rho\,\mathrm{d}x\,\mathrm{d}t,$
we have
$\int_{0}^{T}\int_{B_{R}}\left|\Phi\right|^{2}\rho\,\mathrm{d}x\,\mathrm{d}t\geq\left(1-\dfrac{K}{1+R^{2}}\right)\int_{0}^{T}\int_{\mathbb{R}^{d}}\left|\Phi\right|^{2}\rho\,\mathrm{d}x\,\mathrm{d}t\geq\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left|\Phi\right|^{2}\rho\,\mathrm{d}x\,\mathrm{d}t.$
Therefore,
@@
Next, we claim that when $|x|\leq R$, we have
$|\phi(t,x)|+|\nabla_{x}\phi(t,x)|\leq C\left\|\phi\right\|_{\mu,\alpha}^{\frac{2}{d+5}}.$ (B.104)
Recall that we only care the case when $\left\|\phi\right\|_{\mu,\alpha}\leq 1$. We split this claim into two arguments.
Argument 1. We assume that $\exists(t^{*},x^{*})\in[0,T]\times B_{R}$ s.t. $\left|\phi(t^{*},x^{*})\right|=\varepsilon>0$. Without loss of generality, we assume $\varepsilon\leq 1$. Denote $r:=\varepsilon/(4K)$. Then, for any
$(t,x)\in\widetilde{B}:=\left\{(t,x)\in[0,T]\times\mathbb{R}^{d}~|~\left|(t,x)-(t^{*},x^{*})\right|\leq r\right\},$
we have
$\left|\phi(t,x)\right|\geq\left|\phi(t^{*},x^{*})\right|-\left|\phi(t,x)-\phi(t^{*},x^{*})\right|\geq\varepsilon-2K\cdot\frac{\varepsilon}{4K}=\frac{1}{2}\varepsilon.$
@@
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\left|\phi(t,x)\right|^{2}\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t\geq\int_{\widetilde{B}}\left|\phi(t,x)\right|^{2}\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t\geq\big|\widetilde{B}\big|\,(\frac{1}{2}\varepsilon)^{2}c_{l}\exp(-C_{l}(R^{2}+1))\geq c\,\varepsilon^{d+3}.$
Argument 2. We assume that $\exists(t^{*},x^{*})\in[0,T]\times B_{R}$ s.t. $\left|\nabla_{x}\phi(t^{*},x^{*})\right|=\varepsilon>0$. Without loss of generality, we assume $\varepsilon\leq 1$ and $\left|\nabla_{x}\phi_{1}(t^{*},x^{*})\right|=\varepsilon_{1}>\varepsilon/\sqrt{n}$ (at least one row of $\nabla_{x}\phi$ must be grater than $|\nabla_{x}\phi|/\sqrt{n}$). Let
$v=\dfrac{\nabla_{x}\phi_{1}(t^{*},x^{*})}{\left|\nabla_{x}\phi_{1}(t^{*},x^{*})\right|},\quad\Delta t=r_{1}=\dfrac{\varepsilon_{1}}{16K}\wedge T.$
Then, for any $(t,x)\in[0,T]\times\mathbb{R}^{d}$ s.t. $|t-t^{*}|\leq\Delta t$ and $|x-x^{*}|\leq r_{1}$, we have
$\left|\nabla_{x}\phi_{1}(t,x)-\nabla_{x}\phi_{1}(t^{*},x^{*})\right|\leq 2K(|t-t^{*}|+|x-x^{*}|)\leq\frac{1}{4}\varepsilon_{1},$
which implies
$\partial_{s}\phi_{1}(t,x+sv)\big|_{s=0}=\nabla_{x}\phi_{1}(t,x)^{\top}v\geq\frac{3}{4}\varepsilon_{1}.$
Next, let $z\in v^{\perp}$ and $|z|\leq r_{1}$, then
$\nabla_{x}\phi_{1}(t,x+z+sv)^{\top}v\geq\frac{1}{2}\varepsilon_{1},\quad\forall s\in[-r_{1},r_{1}].$
We define $\psi_{t}(z,s):=\phi_{1}(t,x^{*}+z+sv)$, then $\partial_{s}\psi_{t}(z,s)\geq\frac{1}{2}\varepsilon_{1}$. Taking an integration, we obtain
@@
Let $I=[t^{*}-\Delta t,t^{*}+\Delta t]\cap[0,T]$, then $|I|\geq\Delta t$. We further integrate over $t$ and $z$ and get
@@
where $\omega_{d-1}$ is the volume of unit ball in $\mathbb{R}^{d-1}$. Therefore, by (4.1),
$\quad\int_{0}^{T}\int_{\mathbb{R}^{d}}\left|\phi(t,x)\right|^{2}\rho(r,x)\,\mathrm{d}x\,\mathrm{d}t\geq\int_{0}^{T}\int_{B_{R+1}}\left|\phi(t,x)\right|^{2}\rho(r,x)\,\mathrm{d}x\,\mathrm{d}t$
$\geq\int_{0}^{T}\int_{B_{R+1}}\left|\phi_{1}(t,x)\right|^{2}c_{l}\exp(-C_{l}(R+1)^{2})\,\mathrm{d}x\,\mathrm{d}t\geq c_{l}\exp(-C_{l}(R+1)^{2})\,c\,\varepsilon_{1}^{d+5}\geq c\,\varepsilon^{d+5}.$
We remark that Argument 2 has a similar sprit to a special case of the Gagliardo–Nirenberg interpolation inequality [46], where a small $L^{2}$ norm of $\phi$ implies a small $L^{2}$ norm of $\nabla_{x}\phi$, provided that the higher order derivatives are bounded. We also remark that these two arguments are for small $\varepsilon$ and $\left\|\phi\right\|_{\mu,\alpha}$, which is why we assume $\varepsilon\leq 1$. If $|\phi(t,x)|>1,\,\forall(t,x)\in[0,T]\times B_{R}$ or $|\nabla_{x}\phi(t,x)|>1,\,\forall(t,x)\in[0,T]\times B_{R}$, then we can show that $\left\|\phi\right\|_{\mu,\alpha}$ has a positive lower bound of order $\mathcal{O}(1)$, and (A.29) directly implies (B.95).
Combine Argument 1 and Argument 2, we get if $(t,x)\in[0,T]\times B_{R}$ and $|\phi(t,x)|+|\nabla_{x}\phi(t,x)|=\varepsilon$, then
$\left\|\phi\right\|^{2}_{{\mu,\alpha}}=\int_{0}^{T}\int_{\mathbb{R}^{d}}\left|\phi(t,x)\right|^{2}\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t\geq c\,\varepsilon^{d+5}.$
This implies the claim (B.104) holds.
Finally, we combine all the estimations above and get
@@
where the first three inequalities are because of (B.102), (B.103), and (B.104) respectively. Therefore, (B.95) holds. ∎

### v2

@@ B.3 Superlinear growth lemma
In this section, we prove (B.10). The motivation comes from Lemma A.9, which proves that the optimality gap in value functions has a superlinear (actually quadratic in (A.18)) growth with respect to the optimality gap in controls.
@@
Under the conditions of Theorem 4.4,
$\left\|\nabla_{x}V^{\mu,\alpha}-\nabla_{x}V^{\mu,*}\right\|_{\mu,\alpha}\leq C\left\|\alpha-\alpha^{\mu,*}\right\|_{\mu,\alpha}^{1+\chi},$ (B.90)
where $\chi=\frac{2}{d+5}$.
@@
Fix any $(t,x)\in[0,T]\times\mathbb{R}^{d}$. Let $x_{s}:=X^{\mu,\alpha}_{s}$ denote the state process under $(\mu,\alpha)$, with a given initial condition $x_{t}=x$. Let $\alpha_{s}:=\alpha(s,x_{s})$, $\alpha^{*}_{s}:=\alpha^{\mu,*}(s,x_{s})$, $\phi:=\alpha-\alpha^{\mu,*}$ and $\phi_{s}:=\phi(s,x_{s})$. By (A.29) in Lemma A.13, it suffices to prove the lemma in the case where $\left\|\phi\right\|_{\mu,\alpha}\leq 1$.
By Lemma A.9,
@@
Step 1. Bound each component of the gradient $\left\|\partial_{x_{1}}V^{\mu,\alpha}-\partial_{x_{1}}V^{\mu,*}\right\|_{\mu,\alpha}$. Given $(t,x)$, define the tangent SDE [37] by $y_{s}:=\partial_{x_{1}}X_{s}^{\mu,\alpha}$, i.e., the partial derivative of $x_{s}=X^{\mu,\alpha}_{s}$ with respect to the first component of its initial condition. Denote $b^{\mu,\alpha}(t,x):=b(t,x,\mu_{t},\alpha(t,x))$ and $\sigma^{\mu}(t,x):=\sigma(t,x,\mu_{t})$. Then $y_{s}$ has an initial condition $y_{t}=e_{1}$, where $e_{1}$ denotes the first standard basis of $\mathbb{R}^{d}$, and
$\mathrm{d}y_{s}=\nabla_{x}b^{\mu,\alpha}(t,x_{s})y_{s}\,\mathrm{d}s+\left(\nabla_{x}\sigma^{\mu}(s,x_{s})\cdot y_{s}\right)\mathrm{d}W_{s},$
where $\nabla_{x}\sigma^{\mu}(s,x_{s})\cdot y_{s}:=\sum_{i=1}^{d}\partial_{x_{i}}\sigma^{\mu}(s,x)\,(y_{s})_{i}=:\sigma^{y}_{s}$. By Itô’s formula,
@@
Since both $\left|\nabla_{x}b^{\mu,\alpha}\right|=\left|\nabla_{x}b+\nabla_{\alpha}b\,\nabla_{x}\alpha\right|$ and $\left|\nabla_{x}\sigma^{\mu}\right|$ are bounded, we have
@@
Together with $|y_{t}|=1$, Grönwall’s inequality implies
$\mathbb{E}\left[\left|y_{s}\right|^{2}~\big|~x_{t}=x\right]\leq C,\ \forall 0\leq t\leq s\leq T,\ \forall x\in\mathbb{R}^{d},$ (B.92)
where $C$ is uniform in $t$ and $x$.
Using (B.91) and $y_{s}$, we estimate $\partial_{x_{1}}V^{\mu,\alpha}-\partial_{x_{1}}V^{\mu,*}$. Recall that
$\nabla_{\alpha}^{2}H\left(s,x_{s},\mu_{s},(\alpha^{*}_{s}+v\phi_{s})(s,x_{s}),-\nabla_{x}V^{\mu,*}(s,x_{s})\right)=-\nabla_{\alpha}^{2}f-\sum_{i=1}^{d}\nabla_{\alpha}^{2}b_{i}\,\partial_{x_{i}}V^{\mu,*}(s,x_{s}).$
By the chain rule, taking derivative with respect to $(x_{t})_{1}$ yields (first differentiate with respect to $x_{s}$, then multiply $\partial_{(x_{t})_{1}}x_{s}=y_{s}$)
@@
where we omit dependence on $(s,x_{s},\mu_{s},(\alpha^{*}+v\phi)(s,x_{s}))$ and $(s,x_{s})$ whenever the context is clear. By Assumption 4.1, Assumption 4.3, and the estimation for the value function in Lemma A.12, we obtain
$\left|\partial_{(x_{t})_{1}}\nabla_{\alpha}^{2}H\left(s,x_{s},\mu_{s},(\alpha^{*}_{s}+v\phi_{s})(s,x_{s}),-\nabla_{x}V^{\mu,*}(s,x_{s})\right)\right|\leq C(1+|x_{s}|)|y_{s}|,$ (B.93)
@@
For the term $\phi(s,x_{s})$, we have $\partial_{(x_{t})_{1}}\phi(s,x_{s})=\nabla_{x}\phi(s,x_{s})y_{s}$. Therefore, taking derivative of (B.91) with respect to $x_{1}$ yields
$\partial_{x_{1}}V^{\mu,\alpha}(t,x)-\partial_{x_{1}}V^{\mu,*}(t,x)=-\mathbb{E}\Big[\int_{0}^{T}\int_{0}^{1}\int_{0}^{u}\Big(2\phi_{s}^{\top}\nabla_{\alpha}^{2}H\,\nabla_{x}\phi_{s}\,y_{s}+\phi_{s}^{\top}(\partial_{(x_{t})_{1}}\nabla_{\alpha}^{2}H)\phi_{s}\Big)\mathrm{d}v\,\mathrm{d}u\,\mathrm{d}s\Big].$ (B.95)
Substituting the estimations (B.93), (B.94) into (B.95) yields
$\left|\partial_{x_{1}}V^{\mu,\alpha}(t,x)-\partial_{x_{1}}V^{\mu,*}(t,x)\right|$ $\leq C\,\mathbb{E}\Big[\int_{t}^{T}\int_{0}^{1}\int_{0}^{u}(1+|x_{s}|)|\phi_{s}|\left(|\phi_{s}|+|\nabla_{x}\phi_{s}|\right)|y_{s}|\,\mathrm{d}v\,\mathrm{d}u\,\mathrm{d}s\Big]$ (B.96)
$\leq C\,\mathbb{E}\Big[\int_{t}^{T}(1+|x_{s}|)|\phi_{s}|\left(|\phi_{s}|+|\nabla_{x}\phi_{s}|\right)|y_{s}|\,\mathrm{d}s\Big].$
By consecutive applications of (B.96), Hölder’s inequality, (B.92), Fubini’s theorem and tower property,
$\quad\left\|\partial_{x_{1}}V^{\mu,\alpha}-\partial_{x_{1}}V^{\mu,*}\right\|^{2}_{\mu,\alpha}$
$\leq C\,\mathbb{E}_{x_{t}\sim\rho^{{}^{\mu,\alpha}}_{t}}\Big\{\int_{0}^{T}\mathbb{E}\Big[\int_{t}^{T}(1+|x_{s}|)|\phi_{s}|\left(|\phi_{s}|+|\nabla_{x}\phi_{s}|\right)|y_{s}|\,\mathrm{d}s~\Big|~x_{t}\Big]^{2}\mathrm{d}t\Big\}$
$\leq C\,\mathbb{E}_{x_{t}\sim\rho^{{}^{\mu,\alpha}}_{t}}\Big\{\int_{0}^{T}\mathbb{E}\Big[\int_{t}^{T}(1+|x_{s}|)^{2}|\phi_{s}|^{2}\left(|\phi_{s}|+|\nabla_{x}\phi_{s}|\right)^{2}\,\mathrm{d}s~\Big|~x_{t}\Big]\cdot\mathbb{E}\Big[\int_{t}^{T}|y_{s}|^{2}\,\mathrm{d}s~\Big|~x_{t}\Big]\,\mathrm{d}t\Big\}$
$\leq C\,\mathbb{E}_{x_{t}\sim\rho^{{}^{\mu,\alpha}}_{t}}\Big\{\int_{0}^{T}\mathbb{E}\Big[\int_{t}^{T}(1+|x_{s}|^{2})|\phi_{s}|^{2}\left(|\phi_{s}|+|\nabla_{x}\phi_{s}|\right)^{2}\,\mathrm{d}s~\Big|~x_{t}\Big]\,\mathrm{d}t\Big\}$
@@
Applying the same analysis in each dimension yields
@@
Step 2. We estimate the right-hand side of (B.97). Recall that the density function $\rho$ satisfies the Aronson-type bound (4.1). Fix $R>0$ such that $1+R^{2}\geq 2K$. Denote $B_{R}:=\{x\in\mathbb{R}^{d}:|x|\leq R\}$, $B_{R}^{c}:=\{x\in\mathbb{R}^{d}:|x|>R\}$ and omit dependence on $(t,x)$ whenever the context is clear. We get
$(1+R^{2})\int_{0}^{T}\int_{B_{R}^{c}}\left|\Phi\right|^{2}\rho\,\mathrm{d}x\,\mathrm{d}t\leq\int_{0}^{T}\int_{\mathbb{R}^{d}}(1+|x|^{2})\left|\Phi\right|^{2}\rho\,\mathrm{d}x\,\mathrm{d}t\leq K\int_{0}^{T}\int_{\mathbb{R}^{d}}\left|\Phi\right|^{2}\rho\,\mathrm{d}x\,\mathrm{d}t.$
As a result,
$\int_{0}^{T}\int_{B_{R}}\left|\Phi\right|^{2}\rho\,\mathrm{d}x\,\mathrm{d}t\geq\left(1-\dfrac{K}{1+R^{2}}\right)\int_{0}^{T}\int_{\mathbb{R}^{d}}\left|\Phi\right|^{2}\rho\,\mathrm{d}x\,\mathrm{d}t\geq\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{d}}\left|\Phi\right|^{2}\rho\,\mathrm{d}x\,\mathrm{d}t,$
@@
Next, we claim that:
$|\phi(t,x)|+|\nabla_{x}\phi(t,x)|\leq C\left\|\phi\right\|_{\mu,\alpha}^{\frac{2}{d+5}},\ \forall|x|\leq R.$ (B.99)
Recall that we only have to prove the claim when $\left\|\phi\right\|_{\mu,\alpha}\leq 1$. To proceed, we provide two arguments below.
Argument 1. If there exists $(t^{*},x^{*})\in[0,T]\times B_{R}$, such that $\left|\phi(t^{*},x^{*})\right|=\varepsilon\in(0,1]$ then $\left\|\phi\right\|^{2}_{\rho}\geq c\varepsilon^{d+3}$. Denote $r:=\varepsilon/(4K(R+2))$. For any $(t,x)\in\widetilde{B}$, where
$\widetilde{B}:=\left\{(t,x)\in[0,T]\times\mathbb{R}^{d}:\left|(t,x)-(t^{*},x^{*})\right|\leq r\right\},$
since $\alpha,\alpha^{\mu,*}\in\mathcal{A}$,
$\left|\phi(t,x)\right|$ $\geq\left|\phi(t^{*},x^{*})\right|-\left|\phi(t,x)-\phi(t^{*},x^{*})\right|\geq\varepsilon-(2K(R+1)|t-t^{*}|+2K|x-x^{*}|)$
$\geq\varepsilon-2K(R+2)r=\frac{1}{2}\varepsilon.$
@@
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\left|\phi(t,x)\right|^{2}\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t\geq\iint_{\widetilde{B}}\left|\phi(t,x)\right|^{2}\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t\geq\big|\widetilde{B}\big|\,(\frac{1}{2}\varepsilon)^{2}c_{l}\exp(-C_{l}(R+\tfrac{1}{4K(R+2)})^{2})\geq c\,\varepsilon^{d+3}.$
Argument 2. If there exists $(t^{*},x^{*})\in[0,T]\times B_{R}$, such that $\left|\nabla_{x}\phi(t^{*},x^{*})\right|=\varepsilon\in(0,1]$, then $\left\|\phi\right\|^{2}_{\rho}\geq c\varepsilon^{d+5}$. Note that the norm of at least one row of $\nabla_{x}\phi$ must be greater than $|\nabla_{x}\phi|/\sqrt{n}$, so we assume without loss of generality that $\left|\nabla_{x}\phi_{1}(t^{*},x^{*})\right|=\varepsilon_{1}>\varepsilon/\sqrt{n}$.
Define $v:=\nabla_{x}\phi_{1}(t^{*},x^{*})/\varepsilon_{1},\ r_{1}:=\tfrac{\varepsilon_{1}}{16K},\ \Delta t:=\tfrac{\varepsilon_{1}}{16K}\wedge T.$ For any $(t,x)\in[0,T]\times\mathbb{R}^{d}$ such that $|t-t^{*}|\leq\Delta t$ and $|x-x^{*}|\leq r_{1}$, $\left|\nabla_{x}\phi_{1}(t,x)-\nabla_{x}\phi_{1}(t^{*},x^{*})\right|\leq 2K(|t-t^{*}|+|x-x^{*}|)\leq\frac{1}{4}\varepsilon_{1}$, which implies
$\partial_{s}[\phi_{1}(t,x+sv)]\big|_{s=0}=\nabla_{x}\phi_{1}(t,x)^{\top}v\geq\frac{3}{4}\varepsilon_{1}.$
By Assumption 4.3, $|\nabla_{x}\phi_{1}(t,x+z+sv)-\nabla_{x}\phi_{1}(t,x)|\leq 4Kr_{1},\ \forall z\in v^{\perp},|z|\leq r_{1},s\in[-r_{1},r_{1}]$, implying
$\nabla_{x}\phi_{1}(t,x+z+sv)^{\top}v\geq\frac{1}{2}\varepsilon_{1},\ \forall z\in v^{\perp},|z|\leq r_{1},s\in[-r_{1},r_{1}].$
Define $\psi_{t}(z,s):=\phi_{1}(t,x^{*}+z+sv)$ so that $\partial_{s}\psi_{t}(z,s)\geq\frac{1}{2}\varepsilon_{1}$. Integrating both sides yields $\psi_{t}(z,s)=\psi_{t}(z,0)+\frac{1}{2}\varepsilon_{1}s$. Squaring and integrating both sides once more yield
@@
Set $I=[t^{*}-\Delta t,t^{*}+\Delta t]\cap[0,T]$ so that $|I|\geq\Delta t$. Further integrating with respect to $(t,z)$ yields
@@
where $\omega_{d-1}$ denotes the volume of the unit ball in $\mathbb{R}^{d-1}$. Using (4.1) and $|x|\leq|x^{*}|+|z+sv|\leq R+\frac{1}{8K}$,
$\quad\int_{0}^{T}\int_{\mathbb{R}^{d}}\left|\phi(t,x)\right|^{2}\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t\geq\int_{0}^{T}\int_{B_{R+\frac{1}{8K}}}\left|\phi(t,x)\right|^{2}\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t$
$\geq c_{l}\exp(-C_{l}(R+\tfrac{1}{8K})^{2})\int_{I}\int_{|z|\leq r_{1},z\perp v}\int_{-r_{1}}^{r_{1}}\left|\psi_{t}(z,s)\right|^{2}\mathrm{d}s\,\mathrm{d}z\,\mathrm{d}t\geq c\,\varepsilon_{1}^{d+5}\geq c\,\varepsilon^{d+5}.$
We remark that Argument 2 has a similar sprit to a special case of the Gagliardo–Nirenberg interpolation inequality [46], where a small $L^{2}$ norm of $\phi$ implies a small $L^{2}$ norm of $\nabla_{x}\phi$, provided that the higher order derivatives are bounded. We also remark that, these two arguments require small values of $\varepsilon$. In cases where $|\phi(t,x)|>1,\,\forall(t,x)\in[0,T]\times B_{R}$ or $|\nabla_{x}\phi(t,x)|>1,\,\forall(t,x)\in[0,T]\times B_{R}$, we can always show that $\left\|\phi\right\|_{\mu,\alpha}$ has a positive lower bound of order $\mathcal{O}(1)$, so that (A.29) directly implies (B.90).
Combining Argument 1 and Argument 2: for any $(t,x)\in[0,T]\times B_{R}$ such that $|\phi(t,x)|+|\nabla_{x}\phi(t,x)|=\varepsilon$,
$\left\|\phi\right\|^{2}_{{\mu,\alpha}}=\int_{0}^{T}\int_{\mathbb{R}^{d}}\left|\phi(t,x)\right|^{2}\rho(t,x)\,\mathrm{d}x\,\mathrm{d}t\geq c\,\varepsilon^{d+5},$
which concludes the proof of the claim (B.99).
Finally, combining all previous estimations yields
@@
concluding the proof of (B.90). ∎